OpenStudy (anonymous):

y=9-3^x what is the domain, range and intercepts of this function?

OpenStudy (anonymous):

1) domain is all real numbers because you can raise 3 to any power 2) range is \[(-\infty, 9)\] since \[3^x>0\] and so \[-3^x< 0\] for all x, therefore \[9-3^x<9\] always

OpenStudy (anonymous):

y intercept is what you get when x = 0, so it is \[9-3^0=9-1=8\] or as a coordinate (0,8)

OpenStudy (anonymous):

and finally x intercept is what you get if you put y = 0 solve \[9-3^x=0\] \[9=3^x\] \[x=2\] so it is (2,0)

OpenStudy (anonymous):

Thank you